# Complex Saddles of Charged-AdS Gravitational partition function

### v1

@@ IV KSW allowability of Complex Black Holes
In this section, we study the KSW (Louko–Sorkin–Kontsevich–Segal–Witten) criterion to the generic complex geometry/saddles for the asymptotically AdS spacetimes. We are interested in seeing whether the this could assists in resolving the puzzle pointed in the earlier part of the paper. The KSW is a diagnostic tool to identify “allowable” complex spacetime metrics on which sensible QFTs can be defined Witten:2021nzp ; Louko:1995jw ; Kontsevich:2021dmb . While this seems like a simple requirement, in practice, the KSW puts strong constraints on the metrics which are physically “allowable” Jonas:2022uqb ; Lehners:2023pcn ; Hertog:2023vot ; Hertog:2024nbh . In the following, we examine the compatibility of the complex saddles contributing to the partition function with the KSW criterion. The line-element for the Charged - AdSd+1 spacetime is given by:
@@
with $f(r)$ given the relation (48). Let us consider a generic complex curve $r(u)$ in complex $r$ plane which is monotonically increases along real axis, parameterized by a real parameter $u$ such that $r(0)=r_{+}$ and $r(1)=\infty$. In terms of the $u$- variable, the line element reads as:
@@
Figure 22: A similar plot to fig.(21) corresponding to various regions of grand canonical AdS5 case. The purple saddle in (c) corresponds to negative naked singular geometry. Following values of $\Phi=a\Phi_{\rm max}$ are utilized in plotting: a) $a=1/2$, b) $a=1$ and $a=2$.
@@
Figure 23: A similar plot to figs.(21) and (22) corresponding to various regions of canonical AdS4 case. The purple saddle in all plots correspond to negative naked singular geometry. Following values of $q=aq_{c}$ and $\beta=b\beta_{c}$ are utilized in plotting: a) $a=2,b=2$, b) $a=1,b=1$ and $a=1/2,b=1/2$.
@@ V Conclusion
In this paper, we explore the issues raised recently in the work Mahajan:2025bzo where the authors focused their attention on analysing subtleties of the Euclidean gravitational AdS partition function. They noticed that such partition function in the saddle-point approximation can be reduced to one-dimensional integrals which can be analysed using complex analysis and Picard-Lefschetz theory. An interesting puzzle that was seen was the occurrence of complex saddles, which can become dominant in the partition function and are naively expected to govern its behaviour. However, if that is the case, then this would be in tension with AdS/CFT, which is a serious concern. The authors in Mahajan:2025bzo then used Picard-Lefschetz methods to realize that such complex saddles which although become dominant in certain cases and are in conflict with AdS/CFT, actually either become irrelevant or drop out from path-integral under homology-averaging, thereby resolving the conflict.
@@
In this paper, we explore these by considering the Euclidean path-integral for the case of ${\rm AdS}_{d+1}$-Schwarzschild and ${\rm AdS}_{d+1}$-charged spherically symmetries geometries. As Euclidean gravitational path-integrals suffers from conformal factor problem, following the ideas proposed in Marolf:2022ybi , Lorentzian gravitational path-integral is utilized to define a Euclidean one via integral transform, which has been seen to reproduce the results of standard Euclidean-gravitational path-integral at the saddle-points. The thermal partition function is then a one-dimensional integral, which becomes our starting point. This one-dimensional integral is then applied to ${\rm AdS}_{d+1}$-Schwarzschild and ${\rm AdS}_{d+1}$-charged spherically symmetries geometries, where it is analysed using Picard-Lefschetz methods.
@@
We first consider the case of ${\rm AdS}_{d+1}$-Schwarzschild spherically symmetric geometries, which follows from the Einstein’s equations. The thermal partition function $Z(\beta)$ is computed by utilizing the expression of Entropy and Energy of the geometric configuration as a function of $r_{+}$, which is the horizon’s radius. The one-dimensional integral mentioned in eq. (7) over entropy reduces to one-dimensional integral over $r_{+}$ (modulo Jacobians). This one-dimensional integral is studied in the saddle-point approximation. For the case of ${\rm AdS}_{d+1}$-Schwarzschild, there are three saddles: one appearing at $r_{+}=0$ (with degree d3d-3), the other two follow from quadratic equation and are denoted by $r_{+}^{\pm}$. The Kretschmann scalar shows that all the saddle geometries associated with non-zero $r_{+}^{\pm}$ are singular at $r=0$. It is seen that for all $\beta$, the saddle with $r_{+}=0$ correspond to thermal-AdS. For $\beta>\beta_{\rm max}$ (low-temperature), the two saddles $r_{+}^{\pm}$ form a complex-conjugate pair and correspond to naked-singular geometries. For $\beta<\beta_{\rm max}$ (high-temperature), $r_{+}^{\pm}$ are real saddles with $r_{+}^{+}$ corresponding to larger black hole geometry and $r_{+}^{-}$ corresponding to smaller black hole.
@@
We then ask the question about the contribution of such geometries in the gravitational partition function. A naive expectation would be that all saddles will contribute with the dominance coming from those saddles with the hightest $Re({\cal I}(r_{+}))$. If this is the case then one gets into conflict with AdS/CFT where it well known that thermal-AdS reproduces results consistent with the dual CFT in the low temperature regime Witten:1998zw ; Witten:1998qj . This puzzle was tackled in a recent work for ${\rm AdS}_{5}$ Mahajan:2025bzo , where the authors realized that Picard-Lefschetz methods when utilized in evaluating the one-dimensional integral stated in eq. (7), helps in picking only the “relevant saddles”. For the case of ${\rm AdS}_{d+1}$-Schwarzschild, complex saddles corresponding to naked singular geometries which become dominant for certain values of $\beta>\beta_{\rm max}$ are irrelevant in the path-integral. Small sized black holes which are although relevant but sub-dominating for $\beta<\beta_{\rm max}$ drops out of the path-integral after homology averaging. These sub-dominant small-sized black holes also have negative specific heat. It is seen that methods of Picard-Lefschetz systematically picks the relevant saddles by deforming the original integration contour to pass along the Lefschetz thimbles of the relevant saddles.
@@
We also showed analytically that complex saddles corresponding to naked singularities for ${\rm AdS}_{d+1}$-Schwarzschild always are irrelevant. This is easy to prove as the $Im({\cal I}(r_{+}))$ is constant along the steepest ascent and descent contours. As $Im({\cal I}(r_{+}))$ at these saddles is always non-zero, this implies that steepest ascent contours from these saddles will never intersect the original integration contour lying along the positive real-line, where the $Im({\cal I}(r_{+}))$ is zero. This analytically proves that naked-singularities which although are solution to Einstein’s equation and can be dominating in path-integral for certain values of $\beta>\beta_{\rm max}$, are not relevant saddles.
@@
Small-sized black hole saddles for the case of ${\rm AdS}_{d+1}$-Schwarzschild are always relevant, as the $Im({\cal I}(r_{+}))=0$. However, making an unambiguous decision regarding their relevance is not easy, as the flow lines from all the saddles overlap with each other. This is due to a symmetry mentioned in eq. (107) present in the system. Only after the symmetry is broken and degeneracy is lifted, then a clarity regarding the relevance of saddle is achieved. Complex deformations of Newton’s constant $G=|G|e^{i\epsilon}$ is one way to lift the degeneracies, a process first introduced in Ailiga:2025fny . Once the degeneracies are lifted, it is seen that steepest ascent contours from the will always intersect the original integration contour. These small BHs which have negative specific heat are shown to be always sub-dominating in the path-integral. The process of lifting degeneracies via complex deformation of $G$ however show that the orientation of thimbles in the complex plane changes depending on sign of $\epsilon$. Small-BHs saddles undergo homology jumps as the sign of $\epsilon$ flips. The contribution from these sub-dominant saddles having negative specific heat drops out from path-integral in homology-averaging.
@@
We then studied the case of charged ${\rm AdS}_{d+1}$ black holes probing the same issues investigated for ${\rm AdS}_{d+1}$-Schwarzschild BHs. We first focus on the grand canonical case. Here, we observe that complex saddles and saddles with negative real $r_{+}$ correspond to naked singularities. As before, they are dominating in regimes where $\Phi<\Phi_{\rm max}$ and $\beta>\beta_{\rm max}$ (regime of complex saddle), and for $\Phi>\Phi_{\rm max}$ (regime of negative real saddle). This is the low temperature regime. As before, if the path-integral is performed naively, by taking into account the contribution from all saddles, then we will get into conflict with the AdS/CFT results. However, when situation is analyzed using Picard-Lefschetz, then it is seen these saddles don’t contribute in the partition function as they are either irrelevant (which is the case for complex saddles), or they undergo homology jumps thereby dropping out under homology averaging. For low temperature, the relevant saddles are either thermal-AdS or large-BH, which are the only ones contributing to the partition function thereby resolving the conflicts.
@@
We then discuss the high-temperature phase for the charged-AdS in a similar way. Here also we notice that in the various regimes of the parameters the saddles are real (there are no complex saddles in this regime). These real saddles as before have $Im({\cal I}(r_{+}))=0$, thereby implying the steepest ascent from them intersect the integration contour. However, as all the saddles satisfy this criterion, there is overlapping of flow-lines making it impossible to decide upon the relevance of each saddle unambiguously. Here again we need to employ complex deformation of Newton’s constant $G=|G|e^{i\epsilon}$ in order to lift the degeneracy. All the saddles are relevant once the degeneracy is resolved. Depending on the sign of $\epsilon$, the thimbles structure in the complex plane undergo a homology jump in their pattern. The sub-dominating saddle at $r_{+}^{-}$ having negative specific heat drops out of the partition function on homology averaging despite being a relevant saddle.
@@
Moving on to canonical ensemble for the charged-AdS, we notice a similar pattern. However, due to algebraic limitations, we could only do a study of ${\rm AdS}_{4,\,5,\,6}$. In ${\rm AdS}_{4}$ case we analyse the low-temperature and high-temperature regime, but for ${\rm AdS}_{5,\,6}$ only low-temperature analysis was possible analytically, beside investigating the regimes based on critical charge $q_{c}$ (see table 3 for various regimes). Here again we realize that saddles corresponding to naked-singular geometries (which are either complex saddles or saddles with $r_{+}<0$) don’t contribute in the partition function despite being either dominant saddles in these regimes. They drop out either because they are irrelevant. For high temperature, sub-dominant saddles are relevant. However, they drop out of the partition function under homology averaging. Only thermal-AdS and/or saddles with positive specific heat contribute to the partition function
@@
In this paper we studied the Euclidean partition function for various $\rm AdS$} geometries in arbitrary dimensions and noticed a pattern. Naked singular geometries (which correspond to either complex saddles or saddles with negative $r_{+}$) despite being solution to Einstein’s equation and could be dominant saddles in certain regimes of parameter space, drop out of the partition function. Based on Picard-Lefschetz analysis it is seen that they are irrelevant. Sub-dominant saddles with positive $r_{+}$ come with negative specific heat. Although being relevant in the Picard-Lefschetz analysis, they drop out under homology averaging thereby not contributing to the partition function. Only thermal-AdS saddle and stable-BH are relevant and contribute to the partition function for the cases investigated in this paper.
@@
We also investigate how these saddles fare in light of the KSW allowability criterion. We make use of a weaker version of KSW to study saddles allowability. It is seen that complex saddles which correspond to naked singularities are always disallowed for large enough $\beta$. Meanwhile, for $\beta$ slightly above $\beta_{\rm max}^{\phi}$, weak KSW allows these complex saddles. We believe a more sophisticated version of KSW will make all such complex saddles corresponding to naked singularity disallowed. Naked singularities corresponding to saddles with negative real $r_{+}<0$ are KSW allowed. Sub-dominant small sized black holes with $r_{+}>0$ are also KSW allowed, despite dropping out from the partition function. Thermal-AdS and large-sized black holes are KSW allowed. We believe that this allowability of real saddles corresponding to naked-singularity with negative $r_{+}<0$ and small-sized BHs with $r_{+}>0$ is an anomaly and suggesting that KSW criterion is an insufficient to rule out such “un-physical” geometries. Such situations have been witnessed earlier Witten:2021nzp , where it has been noted that there are geometries which are un-physical but are allowed by KSW criterion. Perhaps a upgraded version of KSW criterion is needed to rule-out such configurations.
@@ Acknowledgments
We are thankful to Justin David and Chethan Krishnan for helpful discussions at various stages of the work. We also like thank Pradipta Pathak for discussion related to BHs.

### v2

@@ IV KSW allowability of Complex Black Holes
In this section, we study the KSW (Louko–Sorkin–Kontsevich–Segal–Witten) allowability of generic complex geometry/saddles for the asymptotically AdS spacetimes. We are interested in seeing whether this could help in resolving the puzzle pointed out in the earlier part of the paper. The KSW is a diagnostic tool to identify “allowable” complex spacetime metrics on which sensible QFTs can be defined Witten:2021nzp ; Louko:1995jw ; Kontsevich:2021dmb . While this seems like a simple requirement, in practice, the KSW puts strong constraints on the metrics which are physically “allowable” Jonas:2022uqb ; Lehners:2023pcn ; Hertog:2023vot ; Hertog:2024nbh ; Liu:2024zti . In the following, we examine the compatibility of the complex saddles contributing to the partition function with the KSW criterion. The line-element for the Charged-AdSd+1 spacetime is given by:
@@
with $f(r)$ given the relation (48). Let us consider a generic complex curve $r(u)$ in complex $r$ plane which monotonically increases along real axis, parameterized by a real parameter $u$ such that $r(0)=r_{+}$ and $r(1)=\infty$. In terms of the $u$- variable, the line element reads as:
@@
Figure 23: A similar plot to fig.(22) corresponding to various regions of grand canonical AdS5 case. The purple saddle in (c) corresponds to negative naked singular geometry. Following values of $\Phi=a\Phi_{\rm max}$ are utilized in plotting: a) $a=1/2$, b) $a=1$ and c) $a=2$.
@@
Figure 24: A similar plot to figs.(22) and (23) corresponding to various regions of canonical AdS4 case. The purple saddle in all plots correspond to negative naked singular geometry. Following values of $q=aq_{c}$ and $\beta=b\beta_{c}$ are utilized in plotting: a) $a=2,b=2$, b) $a=1,b=1$ and c) $a=1/2,b=1/2$.
@@ V Conclusion
In this paper, we explore the issues raised recently in the work Mahajan:2025bzo where the authors focused their attention on analysing subtleties of the Euclidean gravitational AdS partition function. They noticed that such partition function in the saddle-point approximation can be reduced to one-dimensional integrals which can be studied using complex analysis and Picard-Lefschetz theory. An interesting puzzle that was seen is the occurrence of complex saddles, which can become dominant in the partition function and are naively expected to govern its behaviour. However, if that is the case, then this would be in tension with AdS/CFT, which is a serious concern. The authors in Mahajan:2025bzo then used Picard-Lefschetz methods to realize that such complex saddles which although become dominant in certain cases and are in conflict with AdS/CFT, actually either become irrelevant or drop out from path-integral under homology-averaging, thereby resolving the conflict.
@@
In this paper, we explore these features, by considering the Euclidean partition function for the case of ${\rm AdS}_{d+1}$-Schwarzschild and ${\rm AdS}_{d+1}$-charged spherically symmetric geometries. As Euclidean gravitational path-integral suffers from the conformal factor problem, following the ideas proposed in Marolf:2022ybi , Lorentzian gravitational path-integral is utilized to define a Euclidean one via an integral transform, which has been seen to reproduce the results of standard Euclidean-gravitational path-integral at the saddle-points. The thermal partition function is then a finite-dimensional integral, which becomes our starting point. These one-or two-dimensional integral is then applied to ${\rm AdS}_{d+1}$-Schwarzschild and charged ${\rm AdS}_{d+1}$ geometries, respectively, where they are analysed using Picard-Lefschetz methods.
@@
We first consider the case of ${\rm AdS}_{d+1}$-Schwarzschild geometries, which follows from Einstein’s equation. The thermal partition function $Z(\beta)$ is computed by utilizing the expression of Entropy and Energy as a function of $r_{+}$, which is the horizon’s radius. The one-dimensional integral mentioned in eq. (7) over entropy reduces to one-dimensional integral over $r_{+}$ (modulo Jacobians). This one-dimensional integral is studied in the saddle-point approximation. For the case of ${\rm AdS}_{d+1}$-Schwarzschild, there are three saddles: one appearing at $r_{+}=0$ (with degree d3d-3), the other two follow from quadratic equation and are denoted by $r_{+}^{\pm}$. The Kretschmann scalar shows that all the saddle geometries associated with non-zero $r_{+}^{\pm}$ are singular at $r=0$. It is seen that for all $\beta$, the saddle with $r_{+}=0$ corresponds to thermal-AdS. For $\beta>\beta_{\rm max}$ (low-temperature), the two saddles $r_{+}^{\pm}$ form a complex-conjugate pair and correspond to naked-singular geometries. For $\beta<\beta_{\rm max}$ (high-temperature), $r_{+}^{\pm}$ are real saddles with $r_{+}^{+}$ corresponding to larger black hole geometry and $r_{+}^{-}$ corresponding to smaller black hole.
@@
We then ask the question about the contribution of such geometries in the gravitational partition function. A naive expectation would be that all saddles will contribute with the dominance coming from those saddles with the hightest $\rm Re({\cal I}(r_{+}))$. If this is the case then one gets into conflict with AdS/CFT where it well known that thermal-AdS reproduces results consistent with the dual CFT in the low temperature regime Witten:1998zw ; Witten:1998qj . This puzzle was tackled in a recent work for ${\rm AdS}_{5}$ Mahajan:2025bzo , where the authors realized that Picard-Lefschetz method when utilized in evaluating the one-dimensional integral stated in eq. (7), helps in picking only the “relevant saddles” which might not be the dominant saddles. For the case of ${\rm AdS}_{d+1}$-Schwarzschild, complex saddles corresponding to naked singular geometries which become dominant for certain values of $\beta>\beta_{\rm max}$ (see fig. 1) are irrelevant in the partition function. Small black holes which are although relevant but sub-dominating for $\beta<\beta_{\rm max}$ drops out of the path-integral after homology averaging. These sub-dominant small-sized black holes also have negative specific heat. It is seen that methods of Picard-Lefschetz systematically picks the relevant saddles by deforming the original integration contour to pass along the Lefschetz thimbles of the relevant saddles.
@@
We also showed analytically that complex saddles corresponding to naked singularities for ${\rm AdS}_{d+1}$-Schwarzschild are always irrelevant. This is easy to see by noting that the $\rm Im({\cal I}(r_{+}))$ is conserved along the steepest ascent and descent contours. As $\rm Im({\cal I}(r_{+}))$ at these saddles is always non-zero, steepest ascent contours from these saddles will never intersect the original integration contour lying along the positive real line, where the $\rm Im({\cal I}(r_{+}))$ is zero. This analytically proves that naked-singularities, which are solutions to Einstein’s equation and can be dominating in the partition function for certain values of $\beta>\beta_{\rm max}$, are not relevant saddles.
@@
Small-sized black hole saddles for the case of ${\rm AdS}_{d+1}$-Schwarzschild are always relevant, as the $\rm Im({\cal I}(r_{+}))=0$. However, making an unambiguous decision regarding their relevance is not easy, as the flow lines from all the saddles overlap with each other. This is due to a symmetry mentioned in eq. (108) present in the system. Only after the symmetry is broken and degeneracy is lifted, the relevance of saddles can be determined. Complex deformation of Newton’s constant $G=|G|e^{i\epsilon}$ is one way to lift the degeneracies, a process first introduced in Ailiga:2025fny . Once the degeneracies are lifted, it is observed that the steepest ascent contours from all the saddles intersect the original integration contour, and hence, they all contribute to the partition function. The process of lifting degeneracies via complex deformation of $G$, however, shows that the orientation of thimbles in the complex plane changes depending on the sign of $\epsilon$. The descent thimble emanating from Small-BH saddle undergoes homology jumps as the sign of $\epsilon$ flips. The contribution from these sub-dominant saddles having negative specific heat drops out from the partition function in homology-averaging.
@@
We then studied the case of charged ${\rm AdS}_{d+1}$ black holes, addressing the same issues investigated for ${\rm AdS}_{d+1}$-Schwarzschild BHs. We first focus on the grand canonical case. Here, we observe that complex saddles and saddles with negative real $r_{+}$ correspond to naked singularities. As before, they are dominating in regimes where $\beta>\beta_{\rm max}$ and $\Phi<\Phi_{\rm max}$ (regime of complex saddle), and for $\Phi>\Phi_{\rm max}$ (regime of negative real saddle). As before, if the partition function is computed naively at low temperature, by taking into account the contributions from all saddles, we will run into conflict with the AdS/CFT results. However, when the situation is analysed using Picard-Lefschetz, then it is seen that these saddles corresponding to naked singular geometries don’t contribute to the partition function as they are irrelevant (which is the case for negative/complex saddles). For low temperature, the relevant saddles are either thermal-AdS or real positive BH, which are the only ones contributing to the partition function thereby resolving the conflicts.
@@
We then discuss the high-temperature phase for the charged AdS in a similar way. Here, we also notice that in the various regimes of potential ($\Phi$), the saddles are real (there are no complex saddles in this regime). These real saddles as before have $\rm Im({\cal I}(r_{+}))=0$, thereby implying the steepest ascent from them intersect the integration contour. However, as all the saddles satisfy this criterion, there is overlapping of flow-lines making it impossible to decide upon the relevance of each saddle unambiguously. Here again we need to employ complex deformation of Newton’s constant $G=|G|e^{i\epsilon}$ in order to lift the degeneracy. All the positive real saddles are relevant once the degeneracy is resolved. Depending on the sign of $\epsilon$, the thimbles structure in the complex plane undergo a homology jump. The sub-dominating saddle at $r_{+}^{-}$ having negative specific heat drops out of the partition function on homology averaging despite being a relevant saddle.
@@
Moving on to the canonical ensemble for the charged AdS, we notice a similar pattern. However, due to algebraic limitations, we could only do a study of ${\rm AdS}_{4,\,5,\,6}$. In ${\rm AdS}_{4}$ case, we analyse for arbitrary temperature (see table 3 for various regimes), but for ${\rm AdS}_{5,\,6}$ only low-temperature analysis was possible analytically. However, one can argue that the same will hold for arbitrary dimensions. Here again we realize that saddles corresponding to naked-singular geometries (which are either complex saddles or saddles with $r_{+}<0$) don’t contribute to the partition function despite being dominant saddles in low-temperature regime. They are irrelevant saddles according to PL method. For high temperature, sub-dominant saddles (intermediate black holes) are relevant. However, they drop out of the partition function under homology averaging. Only thermal-AdS and/or saddles with positive specific heat contribute to the partition function.
@@
In this paper, we studied the Euclidean partition function for various $\rm AdS$} geometries in arbitrary dimensions and noticed a pattern. Naked singular geometries (which correspond to either complex saddles or saddles with negative $r_{+}$) despite being solution to Einstein’s equation and could be dominant saddles in certain regimes of parameter space, drop out of the partition function. Based on Picard-Lefschetz analysis it is seen that they are irrelevant. Sub-dominant saddles with positive $r_{+}$ come with negative specific heat. Although being relevant in the Picard-Lefschetz analysis, they drop out under homology averaging thereby not contributing to the partition function. Only thermal-AdS saddle and stable-BH are relevant and contribute to the partition function for the cases investigated in this paper.
@@
We also investigate how these saddles fare in light of the KSW allowability criterion. We make use of a weaker version of KSW to study saddles allowability. It is seen that complex saddles which correspond to naked singularities are always disallowed for large enough $\beta$. Meanwhile, for $\beta$ slightly above $\beta_{\rm max}^{\phi}$, weak KSW allows these complex saddles. We believe a more sophisticated version of KSW will make all such complex saddles corresponding to naked singularities disallowed. Naked singularities corresponding to saddles with negative real $r_{+}<0$ are KSW allowed. Sub-dominant small sized black holes with $r_{+}>0$ are also KSW allowed, despite dropping out from the partition function. Thermal-AdS and large-sized black holes are KSW allowed. We believe that this allowability of real saddles corresponding to naked-singularity with negative $r_{+}<0$ and small-sized BHs with $r_{+}>0$ is an anomaly and suggesting that KSW criterion is insufficient to rule out such “unphysical” geometries. Such situations have been witnessed earlier Witten:2021nzp , where it has been noted that there are geometries which are unphysical but are allowed by KSW criterion. Perhaps an upgraded version of the KSW criterion is needed to rule out such configurations.
An interesting future direction will be to extend the analysis to the complex chemical potential $(\Phi)$ case, which arises in the context of gravitational index computation and plays a central role in counting black-hole microstates Jones:2025gno . It would be worthwhile to explore the relevance of such complex solutions utilizing Picard-Lefschetz theory. This extension, however, is expected to be non-trivial, as gravitational action becomes genuinely complex; consequently, the arguments presented in this paper need suitable modifications.
@@ Acknowledgments
We are thankful to Justin David and Chethan Krishnan for helpful discussions at various stages of the work. We also like to thank Pradipta Pathak for the discussion related to BHs. GN would like to acknowledge the startup support from IISc which allowed purchase of workstation where some of the simulations were performed.